OpenStudy (anonymous):

Find dy / dx by implicit differentiation. 9 (x^2)y + 8 (y^2)x = -1

OpenStudy (anonymous):

Okay, what is :\[ \frac {d}{dx} 9x^2y \]

OpenStudy (anonymous):

We can start by using product rule: \[ \frac{d}{dx}9x^2y= \left(\frac{d}{dx}9x^2\right)\cdot y + 9x^2\cdot \left(\frac{d}{dx} y\right) \]

OpenStudy (anonymous):

Are you still stuck?

OpenStudy (anonymous):

No i just made my first question and playing around with the website to figure it all out but thanks!

OpenStudy (anonymous):

Oh, okay so you did figure out out then?

OpenStudy (anonymous):

yep :)